шiлili b, y + бу 0; у(0) 0, (0) 1; 4y + 5 -- 2y o Problem 4. Sоlvе hоtоgenous epuations
Answer #1

→ (9) - y - 4y + 4 = 0 4D + +) m2 4m + f = 0 m-2)0 me 2,2 m3-m2-3m2 +3m +2m -2=0 m?cm-1) - 3mcm-1) + 2 cm+) =0 (m2-3m +2j (